# AD-GS: Object-Aware B-Spline Gaussian Splatting for Self-Supervised Autonomous Driving

### v2

@@ 4.1 Setup
Datasets. We evaluate our model using three widely used datasets: KITTI [7], Waymo [29] and nuScenes [1]. For the KITTI dataset, we follow the settings of SUDS [32], which include three sequences, two cameras, and a resolution of $375\times 1242$ while using images of different proportions for training. For the Waymo dataset, we mainly follow the settings of StreetGS [35], which consists of eight sequences and a single camera, but we adopt a higher resolution of $1280\times 1920$. For the nuScenes dataset, we select six sequences with three cameras and 60 frames, setting the resolution to $900\times 1600$. We select one out of every four frames in the nuScenes sequences for testing, while the remaining frames are used for training.
@@ 4.2 Comparison
Comparison with Self-Supervised Models. We compare AD-GS with the state-of-the-art self-supervised models on the KITTI [7], Waymo [29] and nuScenes [1] datasets. The results are presented in Table 1 and Figure 4 for KITTI, Table 2 and Figure 5 for Waymo, and Table 3 and Figure 6 for nuScenes. Notably, all the LPIPS [41] evaluations on the KITTI dataset are based on AlexNet [13], while the evaluations on the other datasets use VGG [28]. Experimental results across all three datasets demonstrate that our model significantly outperforms the state-of-the-art self-supervised models. Compared to PVG [2], which relies solely on trigonometric functions for motion representation, our method benefits from the local fitting properties of learnable B-spline curves, as illustrated in Figure 4. Additionally, as shown in Figures 5 and 6, the object-aware splatting technique in AD-GS, which incorporates the simplified 2D segmentation and the bidirectional temporal visibility masks, achieves higher rendering quality than EmerNeRF [36], which uses feature supervision to help scene decomposition. More results can be found in Section C of our supplementary.
Comparison with Manually Assisted Models. As shown in Table 1 and Table 2, AD-GS achieves comparable performance despite not relying on manual 3D annotations for motion and scene reconstruction. These results highlight the strong performance of AD-GS in the self-supervised scene rendering. However, when the training views become too sparse, such as the setting “KITTI-25%” in Table 1 which only uses 25% images for training, self-supervised models might have difficulties in fitting the correct object motions by themselves. However, our model still achieves improvements compared to self-supervised models under the extremely sparse view setting.
@@ 4.3 Ablation Study
Figure 7: Loss ablation on the KITTI [7] dataset by gradually adding the losses.
Ablation of Losses. We evaluate the impact of each loss term under the “KITTI-75%” setting by progressively adding $\mathcal{L}_{obj}\&\mathcal{L}_{sky}$, $\mathcal{L}_{f}\&L_{d}$ and $\mathcal{L}_{r}$ in Equation 3.4, corresponding to the “obj&sky”, “flow&depth” and “reg”, respectively. As shown in Table 4 and Figure 7, the 2D segmentation guidance provided by $\mathcal{L}_{obj}\&\mathcal{L}_{sky}$ helps the object and background Gaussians adjust their positions, leading to clearer rendering. Additionally, $\mathcal{L}_{d}\&\mathcal{L}_{f}$ contribute extra motion and 3D information, enabling AD-GS to reconstruct the scene more accurately. The regularization term $\mathcal{L}_{r}$ helps AD-GS maintain structural coherence, preventing disorganization and reducing artifacts.
@@
Figure 8: Object modeling module ablation on the Waymo [29] dataset by gradually adding the modules.
@@ 5 Conclusion
In this paper, we introduce AD-GS, a self-supervised Gaussian-based model for high-quality auto-driving scene rendering without relying on the manual 3D annotations. AD-GS utilizes the learnable B-spline curves and B-spline quaternion curves, combined with trigonometric functions, to provide both local and global motion fitting. To achieve more accurate and robust scene modeling, we decompose each scene into objects and background based on the simplified pseudo 2D segmentation results. The objects are represented by dynamic Gaussians with bidirectional temporal visibility masks to handle sudden appearances and disappearances, while the background remains static. In addition to the self-supervision with pseudo labels, we propose the visibility and physically rigid regularization techniques to prevent chaotic behavior in AD-GS. Our experiments demonstrate the state-of-the-art performance achieved by AD-GS, and we hope AD-GS will inspire better works to promote the development of self-supervised rendering models for auto-driving scenes.

### v3

@@ 4.1 Setup
Datasets. We evaluate our model using three widely used datasets: KITTI [7], Waymo [29] and nuScenes [1]. For the KITTI dataset, we follow the settings of SUDS [32] to conduct experiments on three sequences with the resolution of $375\times 1242$ captured by two cameras, where different proportions of images are used for training. For the Waymo dataset, we mainly follow the settings of StreetGS [35], including eight sequences captured by a single camera, but we adopt a higher resolution of $1280\times 1920$. For the nuScenes dataset, we select six sequences with three cameras and 60 frames, setting the resolution to $900\times 1600$. We select one out of every four frames in the nuScenes sequences for testing, while the remaining frames are used for training.
@@ 4.2 Comparison
Comparison with Self-Supervised Models. We compare AD-GS with the state-of-the-art self-supervised models on the KITTI [7], Waymo [29] and nuScenes [1] datasets. The results are presented in Table 1 and Figure 4 for KITTI, Table 2 and Figure 5 for Waymo, and Table 3 and Figure 6 for nuScenes. Notably, all the LPIPS [41] evaluations on the KITTI dataset are based on AlexNet [13], while the evaluations on the other datasets use VGG [28]. Experimental results across all three datasets demonstrate that our model significantly outperforms the state-of-the-art self-supervised models. Compared to PVG [2], which relies solely on trigonometric functions for motion representation, our method benefits from the local fitting properties of learnable B-spline curves, as illustrated in Figure 4. Additionally, as shown in Figures 5 and 6, the object-aware splatting technique in AD-GS, including the simplified 2D segmentation and the bidirectional temporal visibility masks, achieves higher rendering quality than EmerNeRF [36], which uses feature supervision to help scene decomposition. More results can be found in Section C of our supplementary.
Comparison with Manually Assisted Models. As shown in Table 1 and Table 2, AD-GS achieves comparable performance despite not relying on manual 3D annotations for motion and scene reconstruction. These results highlight the good performance of AD-GS in the self-supervised scene rendering. However, when the training views become too sparse, such as the setting “KITTI-25%” in Table 1 which only uses 25% images for training, self-supervised models might have difficulties in fitting the correct object motions by themselves. However, our model still achieves improvements compared to self-supervised models under the extremely sparse view setting.
@@ 4.3 Ablation Study
Figure 7: Visualization of loss ablation on the KITTI [7] dataset by gradually adding the losses.
Ablation of Losses. We evaluate the impact of each loss term under the “KITTI-75%” setting by progressively adding $\mathcal{L}_{obj}\&\mathcal{L}_{sky}$, $\mathcal{L}_{f}\&L_{d}$ and $\mathcal{L}_{r}$ in Equation 3.4, corresponding to the “obj&sky”, “flow&depth” and “reg”, respectively. As shown in Table 4 and Figure 7, the 2D segmentation guidance provided by $\mathcal{L}_{obj}\&\mathcal{L}_{sky}$ helps the object and background Gaussians adjust their positions, leading to clearer rendering. Additionally, $\mathcal{L}_{d}\&\mathcal{L}_{f}$ contribute extra motion and 3D information, enabling AD-GS to reconstruct the scene more accurately. The regularization term $\mathcal{L}_{r}$ encourages AD-GS to maintain structural coherence, preventing disorganization and reducing artifacts.
@@
Figure 8: Visualization of object modeling module ablation on the Waymo [29] dataset by gradually adding the modules.
@@ 5 Conclusion
In this paper, we introduced AD-GS, a self-supervised Gaussian-based model for high-quality auto-driving scene rendering without relying on the manual 3D annotations. AD-GS utilizes the learnable B-spline curves and B-spline quaternion curves, combined with trigonometric functions, to provide both local and global motion fitting. To achieve more accurate and robust scene modeling, we decomposed each scene into objects and background based on the simplified pseudo 2D segmentation results. The objects are represented by dynamic Gaussians with bidirectional temporal visibility masks to handle sudden appearances and disappearances, while the background remains static. In addition to the self-supervision with pseudo labels, we proposed the visibility and physically rigid regularization techniques to prevent chaotic behavior in AD-GS. Our experiments demonstrate the state-of-the-art performance achieved by AD-GS under self-supervised conditions. However, achieving better rendering quality than annotation-dependent models still needs further research.